# MINIMIZATION OF NONSMOOTH WEAKLY CONVEX FUNCTION OVER PROX-REGULAR SET FOR ROBUST LOW-RANK MATRIX RECOVERY

### v1

@@ 1 Introduction
For Problem 1.1, the convex $\ell_{1}$-norm has been utilized as standard loss to detect outliers [10, 1, 2, 3]. For example, an optimization model with an expression of $\mathcal{L}_{r}=\{\bm{U}\bm{V}^{\mathsf{T}}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\bm{U}\in\mathbb{R}^{n_{1}\times r},\bm{V}\in\mathbb{R}^{n_{2}\times r}\}$ and $\lambda\geq 0$:
@@
is expected to improve the estimation performance for Problem 1.1. However, for the model (4), any optimization algorithm has not been reported so far mainly because the cost function in (4) with a weakly convex $\ell$ is not necessarily weakly convex.
@@
    $g:\mathcal{Z}\to\mathbb{R}$ is (a) Lipschitz continuous (possibly nonsmooth), (b) $\eta$-weakly convex with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex, and (c) prox-friendly, i.e., $\mathrm{prox}_{\mu g}\ (\mu\in(0,\eta^{-1}))$ (see (9)) is available as a computable tool;
@@
    $C\subset\mathcal{X}$ is a nonempty closed prox-regular set [^1] Other than $\mathfrak{L}_{r,\sigma}$ in (6), prox-regular sets include, e.g., closed convex sets, $C^{2}$ embedded submanifolds in $\mathcal{X}$ (see, e.g., [23], and [24, Lemma 2.1]), and $\{\bm{x}\in\mathbb{R}^{n}\mid\left\lVert\bm{x}\right\rVert=1,[\bm{x}]_{i}\geq 0\ \forall i\}$ (see [25, Lemma 3 (ii)]).  [21, Thm. 1.3], i.e., the metric projection $P_{C}:\mathcal{X}\rightrightarrows C:\widebar{\bm{x}}\mapsto\mathop{\mathrm{argmin}}_{\bm{x}\in C}\left\lVert\widebar{\bm{x}}-\bm{x}\right\rVert$ onto $C$ is single-valued on some open subset containing $C$. Moreover, we assume that at least one point $\bm{x}\in P_{C}(\widebar{\bm{x}})$ can be computed for every $\widebar{\bm{x}}\in\mathcal{X}$.
@@ 2 Preliminary on Nonsmooth Analysis
The proximity operator and the Moreau envelope have been used as computational tools for nonsmooth optimization [36, 37, 34, 27, 35, 26, 28, 29, 30, 6]. For a Euclidean space $\mathcal{H}$, the proximity operator of a proper function $J:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ with index $\mu>0$ is defined by
@@
$^{\mu}g$ has notable properties as an approximation of the nonsmooth function $g$: (i) $\lim_{\mu\to 0}{}^{\mu}g(\widebar{\bm{z}})=g(\widebar{\bm{z}})\ (\widebar{\bm{z}}\in\mathcal{Z})$; (ii) continuous differentiability of $^{\mu}g$ with $\nabla{}^{\mu}g(\widebar{\bm{z}})=\mu^{-1}(\widebar{\bm{z}}-\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$; (iii) Lipschitz continuity of $\nabla{}^{\mu}g$ (see, e.g, [38, Cor. 3.4]). Both $^{\mu}g$ and $\nabla{}^{\mu}g$ are computable if $g$ is prox-friendly (see Rem. 1.3 for such $g$). Moreover, $\nabla({}^{\mu}g\circ\mathfrak{S})$ enjoys the Lipschitz continuity.
Fact 2.2 (Lipschitzian of $\nabla({}^{\mu}g\circ\mathfrak{S})$ [26, Prop. 4.5(a)]).
@@ 3 Projected variable smoothing algorithm
Fact 3.2 (Sufficient decrease property, e.g., [39, Lemma 2]).
@@
$\ J(\bm{x})\leq J(\widebar{\bm{x}})-c\gamma\left\lVert\frac{\widebar{\bm{x}}-\bm{x}}{\gamma}\right\rVert^{2}\overset{(\clubsuit)}{\leq}J(\widebar{\bm{x}})-c\gamma\left(\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})\right)^{2},$ (13)
where we used $\left\lVert(\widebar{\bm{x}}-\bm{x})/\gamma\right\rVert\geq\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})$ (see (11)) in $(\clubsuit)$.
@@
$\hskip-16.00008pt\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\geq\begin{cases}\mathcal{M}_{\widebar{\gamma}}^{F,\iota_{C}}(\widebar{\bm{x}}),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}>0;\\ \mathop{\mathrm{dist}}\left(\bm{0},\partial_{F}(F+\iota_{C})(\widebar{\bm{x}})\right),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}=0\end{cases}\hskip-10.00002pt$ (14)
@@
In Alg. 1, we update estimates as $\bm{x}_{n+1}\in P_{C}(\bm{x}_{n}-\gamma_{n}\nabla F_{n}(\bm{x}_{n}))$ with a stepsize $\gamma_{n}$ such that the sufficient decrease condition in (13) with $\widebar{\bm{x}}\coloneqq\bm{x}_{n}$ and $J\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$ is achieved by $(\bm{x},\gamma)\coloneqq(\bm{x}_{n+1},\gamma_{n})$, where $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$ with $\alpha\geq 1$ (Note: $\alpha=3$ is recommended in the literature of variable smoothing-type algorithms [34, 35, 26, 27, 28, 6, 30] that can not be applied to Problem 1.2 due to the nonconvexity of $C$). To find such a pair $(\bm{x}_{n+1},\gamma_{n})$, we employ a standard backtracking algorithm in line 4-9 of Alg. 1 (see, e.g., [26, 30]). Thanks to Fact 3.2 with Fact 2.2, we can find such $(\bm{x}_{n+1},\gamma_{n})$ in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{(1-2c)}{\widetilde{\gamma}L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}}\right)\right\rceil\right\}$ backtracking steps in line 4-9, where Alg. 1 does not require any knowledge on a constant $L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ in Fact 2.2.
We present a convergence analysis of Alg. 1 in Theorem 3.4444To show Theorem 3.4, we can apply the discussion similar to the proof of [30, Thm. III.6 with $\phi\coloneqq\iota_{C}$] although $\phi$ in [30] is assumed to be convex (Note: $\iota_{C}$ (see after (8)) with $C$ in Problem 1.2 is nonconvex). .
@@
Consider Problem 1.2. Choose $\bm{x}_{1}\in C$, $c\in(0,2^{-1})$, $\widetilde{\gamma}\in\mathbb{R}_{++}$, and $\alpha\geq 1$. For $(\bm{x}_{n})_{n=1}^{\infty}\subset C$ generated by Alg. 1, the following hold with $F_{n}={}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$ and $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$:
@@
  2. (b)
    Choose a subsequence $(\bm{x}_{m(l)})_{l=1}^{\infty}\subset\mathcal{X}$ of $(\bm{x}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\mathcal{M}_{\gamma_{m(l)}}^{F_{m(l)},\iota_{C}}(\bm{x}_{m(l)})=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing. Then, every cluster point $\bm{x}^{\star}\in C$ of $(\bm{x}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.2.
@@
For simplicity, we are focusing on Problem 1.2 in this paper. Nevertheless, Lemma 3.1, and Thms. 3.3 and 3.4 can be extended to a case where $\iota_{C}$ is replaced by a proper lower semicontinuous and prox-regular[18, Def. 13.27] function $\phi:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ such that $\partial_{F}\phi$ is outer semicontinuous [18, Def. 5.4] at every $\widebar{\bm{x}}\in\mathrm{dom}(\phi)$ (Note: such a $\phi$ includes, e.g., the $\ell_{0}$-pseudonorm). In this case, the measure in (11) is extended to $\mathcal{M}_{\gamma}^{F,\phi}(\widebar{\bm{x}})\coloneqq\inf\left\{\left\lVert\frac{\widebar{\bm{x}}-\bm{p}}{\gamma}\right\rVert\mid\bm{p}\in\mathrm{prox}_{\gamma\phi}(\widebar{\bm{x}}-\gamma\bm{v}),\ \bm{v}\in\partial_{F}F(\widebar{\bm{x}})\right\}$. To the problem for finding a stationary point of $F+\phi$, we can apply a modified version of Alg. 1, in the sense of Thm. 3.4, by replacing (i) $P_{C}$ in line 5 with $\mathrm{prox}_{\rho^{m}\widetilde{\gamma}}$; and (ii) the condition in line 6 with the condition $(F_{n}+\phi)(\bm{x})\leq(F_{n}+\phi)(\bm{x}_{n})-c\rho^{m}\widetilde{\gamma}\left\lVert\frac{\bm{x}_{n}-\bm{x}}{\rho^{m}\widetilde{\gamma}}\right\rVert$, where $\widetilde{\gamma}\in\mathbb{R}_{++}$ is chosen such that $\phi+\frac{1}{2\widetilde{\gamma}}\left\lVert\cdot\right\rVert^{2}$ is bounded below.
@@ 4 Numerical Experiments on Synthetic Data
which is $(\theta-1)^{-1}$-weakly convex (e.g., [34]). For comparisons, we employed two state-of-the-art models: the model (3) in [1] and the model in [3] with weights $\lambda,\beta\in\mathbb{R}_{++}$:
@@
Table 1 demonstrates the averaged RMSE and runtime of each model [^5] For (5) with $\ell\coloneqq r_{\theta}^{\rm SCAD}$ in (15), we employed $\theta\in\{2.5,2.7,2.9,3.1\}$. For (3), $\lambda\in\{10^{-2},10^{-1},1,2,5\}$, was chosen. Moreover, stepsize of the subgradient method used in [1] at $n$th iteration was chosen from $0.95^{n}$ or $2\times 0.95^{n}$, where the first choice was suggested by [1] while the second choice achieved better results in some cases in our scenario. For (16), $\lambda\coloneqq t\sqrt{mn_{2}\log(n_{1}+n_{2})}$ with $t\in\{0.1,0,5\}$ and $\beta\in\{0.1,0.5,0.9\}$ were chosen according to [3]. over $100$ trials with $(n_{1},n_{2},r)=(40,50,5)$, $m=p_{m}n_{1}n_{2}$ and $\lvert\mathcal{I}_{\rm out}\rvert=p_{\lvert\mathcal{I}_{\rm out}\rvert}m$ for $p_{m}\in\{0.3,0.5,0.7\}$ and $p_{\lvert\mathcal{I}_{\rm out}\rvert}\in\{0.7,0.5\}$, where we used RMSE $\lVert\widehat{\bm{X}}-\bm{X}^{\star}\rVert_{F}/\sqrt{n_{1}n_{2}}$ to measure a recovery error with the final estimate $\widehat{\bm{X}}\in\mathbb{R}^{n_{1}\times n_{2}}$ of algorithm. We note that Problem 1.1 becomes challenging as $p_{m}$ is smaller and $p_{\lvert\mathcal{I}_{\rm out}\rvert}$ is larger.
From Table 1, we observe that convergence speeds of the proposed models (5) and the model (16) are much slower than that of the model (3). This is because singular value decompositions (SVD) of $n_{1}$-by-$n_{2}$ matrices are required at every iteration in solvers for (5) (see also Remark 1.3) and for (16) while any SVD is not required in a subgradient method for (3) because of the decomposition $\bm{X}=\bm{U}\bm{V}^{\mathsf{T}}$ in (3) (Note: acceleration of the proposed Alg. 1 is beyond the scope of this paper, and will be addressed in future work). Moreover, from Table 1, the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ achieves smaller RMSE than the model (5) with $\ell\coloneqq\lvert\cdot\rvert$ in a shorter runtime; and RMSE of the existing models (3) and (16) are much higher than that of the model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ in particular for $p_{m}\in\{0.3,0.5\}$. These observations imply that a weakly convex $r^{\rm SCAD}_{\theta}$ is more robust against outliers, at least in the proposed model (5), than a convex $\lvert\cdot\rvert$ as we expected; and the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ outperforms the existing models (3) and (16) in terms of recovery accuracy.

### v2

@@ 1 Introduction
For Problem 1.1, the convex $\ell_{1}$-norm $\left\lVert\cdot\right\rVert_{1}$ has been utilized as standard loss to detect outliers [10, 1, 2, 3]. For example, an optimization model with an expression of $\mathcal{L}_{r}=\{\bm{U}\bm{V}^{\mathsf{T}}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\bm{U}\in\mathbb{R}^{n_{1}\times r},\bm{V}\in\mathbb{R}^{n_{2}\times r}\}$ and $\lambda\geq 0$:
@@
is expected to improve the estimation performance for Problem 1.1. However, for the model (4), any reliable optimization algorithm has not been reported so far mainly because the cost function in (4) with a weakly convex $\ell$ is not necessarily weakly convex.
@@
    $g:\mathcal{Z}\to\mathbb{R}$ is (a) $L_{g}$-Lipschitz continuous with $L_{g}>0$ (possibly nonsmooth), (b) $\eta$-weakly convex with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex, and (c) prox-friendly, i.e., $\mathrm{prox}_{\mu g}\ (\mu\in(0,\eta^{-1}))$ (see (9)) is available as a computable tool;
@@
    $C\subset\mathcal{X}$ is a nonempty closed prox-regular set [^1] Other than $\mathfrak{L}_{r,\sigma}$ in (6), prox-regular sets include, e.g., closed convex sets, $C^{2}$ embedded submanifolds in $\mathcal{X}$ (see, e.g., [23], and [24, Lemma 2.1]), and $\{\bm{x}\in\mathbb{R}^{n}\mid\left\lVert\bm{x}\right\rVert=1,[\bm{x}]_{i}\geq 0\ \forall i\}$ (see [25, Lemma 3 (ii)]).  [21, Thm. 1.3], i.e., the metric projection $P_{C}:\mathcal{X}\rightrightarrows C:\widebar{\bm{x}}\mapsto\mathop{\mathrm{argmin}}_{\bm{x}\in C}\left\lVert\widebar{\bm{x}}-\bm{x}\right\rVert$ onto $C$ is single-valued on some open superset of $C$. Moreover, we assume that at least one point $\bm{x}\in P_{C}(\widebar{\bm{x}})$ can be computed for every $\widebar{\bm{x}}\in\mathcal{X}$.
@@ 2 Preliminary on Nonsmooth Analysis
The proximity operator and the Moreau envelope have been used as computational tools for nonsmooth optimization [36, 34, 27, 35, 26, 28, 29, 30, 6]. For a Euclidean space $\mathcal{H}$, the proximity operator of a proper function $J:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ with index $\mu>0$ is defined by
@@
$^{\mu}g$ has notable properties as an approximation of the nonsmooth function $g$: (i) $\lim_{\mu\to 0}{}^{\mu}g(\widebar{\bm{z}})=g(\widebar{\bm{z}})\ (\widebar{\bm{z}}\in\mathcal{Z})$; (ii) continuous differentiability of $^{\mu}g$ with $\nabla{}^{\mu}g(\widebar{\bm{z}})=\mu^{-1}(\widebar{\bm{z}}-\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$; (iii) Lipschitz continuity of $\nabla{}^{\mu}g$ (see, e.g, [37, Cor. 3.4]). Both $^{\mu}g$ and $\nabla{}^{\mu}g$ are computable if $g$ is prox-friendly (see Rem. 1.3 for such $g$). Moreover, $\nabla({}^{\mu}g\circ\mathfrak{S})$ enjoys the Lipschitz continuity.
Fact 2.2 (Lipschitzian of $\nabla({}^{\mu}g\circ\mathfrak{S})$ [26, Prop. 4.2(a)]).
@@ 3 Projected variable smoothing algorithm
Fact 3.2 (Sufficient decrease property, e.g., [38, Lemma 2]).
@@
$\ J(\bm{x})\leq J(\widebar{\bm{x}})-c\gamma\left\lVert(\widebar{\bm{x}}-\bm{x})/\gamma\right\rVert^{2}\leq J(\widebar{\bm{x}})-c\gamma\left(\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})\right)^{2}.$ (13)
@@
$\hskip-16.00008pt\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\geq\begin{cases}\mathcal{M}_{\widebar{\gamma}}^{F,\iota_{C}}(\widebar{\bm{x}}),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}>0;\\ \mathop{\mathrm{dist}}\left(\bm{0},\partial_{F}(F+\iota_{C})(\widebar{\bm{x}})\right),&\hskip-8.00003pt\par\mathrm{if}\ \widebar{\gamma}=0\end{cases}\hskip-10.00002pt$ (14)
@@
In Alg. 1, we update estimates as $\bm{x}_{n+1}\in P_{C}(\bm{x}_{n}-\gamma_{n}\nabla F_{n}(\bm{x}_{n}))$ with a stepsize $\gamma_{n}$ such that the sufficient decrease condition in (13) with $\widebar{\bm{x}}\coloneqq\bm{x}_{n}$ and $J\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$ is achieved by $(\bm{x},\gamma)\coloneqq(\bm{x}_{n+1},\gamma_{n})$, where $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$ with $\alpha\geq 1$ (Note: $\alpha=3$ is recommended in the literature of variable smoothing-type algorithms [34, 35, 26, 27, 28, 29, 6, 30] that can not be applied to Problem 1.2 due to the nonconvexity of $C$). To find such a pair $(\bm{x}_{n+1},\gamma_{n})$, we employ a standard backtracking algorithm in line 4-9 of Alg. 1 (see, e.g., [26, 30]). Thanks to Fact 3.2 with Fact 2.2, we can find such $(\bm{x}_{n+1},\gamma_{n})$ in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{1-2c}{\widetilde{\gamma}L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}}\right)\right\rceil\right\}$ backtracking steps in line 4-9, where Alg. 1 does not require any knowledge on a constant $L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ in Fact 2.2. Finally, we present a convergence analysis of Alg. 1 in Theorem 3.4.
@@
Consider Problem 1.2. Choose $\bm{x}_{1}\in C$, $c\in(0,2^{-1})$, $\rho\in(0,1)$, $\widetilde{\gamma}\in\mathbb{R}_{++}$, and $\alpha\geq 1$. For $(\bm{x}_{n})_{n=1}^{\infty}\subset C$ generated by Alg. 1, the following hold with $F_{n}={}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$ and $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$:
@@
    If $\alpha>1$, then $\min\limits_{1\leq n\leq k}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\leq\sqrt{\frac{\chi}{(k+1)^{1-\alpha^{-1}}-1}}$} $(k\in\mathbb{N})$ holds with $\chi\coloneqq\frac{\left(F(\bm{x}_{1})-\inf_{\bm{x}\in C}F(\bm{x})+L_{g}^{2}(2\eta)^{-1}\right)(\varpi_{1}+2\eta\varpi_{2})}{2\eta c\min\left\{\widetilde{\gamma}(\varpi_{1}+2\eta\varpi_{2}),\rho(1-2c)\right\}}\in\mathbb{R}_{++}$, where $\varpi_{1},\varpi_{2}\in\mathbb{R}_{++}$ are given in Fact 2.2.
  2. (b)
@@
  3. (c)
    Choose a subsequence $(\bm{x}_{m(l)})_{l=1}^{\infty}\subset\mathcal{X}$ of $(\bm{x}_{n})_{n=1}^{\infty}$ satisfying $\lim_{l\to\infty}\mathcal{M}_{\gamma_{m(l)}}^{F_{m(l)},\iota_{C}}(\bm{x}_{m(l)})=0$. Then, every cluster point $\bm{x}^{\star}\in C$ of $(\bm{x}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.2.
@@
For simplicity, we are focusing on Problem 1.2 in this paper. Nevertheless, Lemma 3.1, and Thms. 3.3 and 3.4 can be extended to a case where $\iota_{C}$ is replaced by a proper lower semicontinuous and prox-regular[18, Def. 13.27] function $\phi:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ such that $\partial_{F}\phi$ is outer semicontinuous [18, Def. 5.4] at every $\widebar{\bm{x}}\in\mathrm{dom}(\phi)$ (Note: such a $\phi$ includes, e.g., the $\ell_{0}$-pseudonorm). In this case, the measure in (11) is extended to $\mathcal{M}_{\gamma}^{F,\phi}(\widebar{\bm{x}})\coloneqq\inf\left\{\left\lVert(\widebar{\bm{x}}-\bm{p})/\gamma\right\rVert\mid\bm{p}\in\mathrm{prox}_{\gamma\phi}(\widebar{\bm{x}}-\gamma\bm{v}),\ \bm{v}\in\partial_{F}F(\widebar{\bm{x}})\right\}$. To the problem for finding a stationary point of $F+\phi$, we can apply a modified version of Alg. 1 by replacing (i) $P_{C}$ in line 5 with $\mathrm{prox}_{\rho^{m}\widetilde{\gamma}}$; and (ii) the condition in line 6 with the condition $(F_{n}+\phi)(\bm{x})\leq(F_{n}+\phi)(\bm{x}_{n})-c\rho^{m}\widetilde{\gamma}\left\lVert(\bm{x}_{n}-\bm{x})/(\rho^{m}\widetilde{\gamma})\right\rVert^{2}$.
@@ 4 Numerical Experiments on Synthetic Data
which is $(\theta-1)^{-1}$-weakly convex (e.g., [34]). For comparisons, we employed two state-of-the-art models: the model (3) in [1] and the following model in [3] with weights $\lambda,\beta\in\mathbb{R}_{++}$:
@@
Table 1 demonstrates the averaged root-mean-square error (RMSE) and runtime of each model [^4] For (5) with $\ell\coloneqq r_{\theta}^{\rm SCAD}$ in (15), we employed $\theta\in\{2.5,2.7,2.9,3.1\}$. For (3), $\lambda\in\{10^{-2},10^{-1},1,2,5\}$ was chosen. Moreover, stepsize of the subgradient method used in [1] at $n$th iteration was chosen from $0.95^{n}$ or $2\times 0.95^{n}$, where the first choice was suggested by [1] while the second choice achieved better results in some cases in our scenario. For (16), $\lambda\coloneqq t\sqrt{mn_{2}\log(n_{1}+n_{2})}$ with $t\in\{0.1,0,5\}$ and $\beta\in\{0.1,0.5,0.9\}$ were chosen according to [3]. over $100$ trials with $(n_{1},n_{2},r)=(40,50,5)$, $m=p_{m}n_{1}n_{2}$ and $\lvert\mathcal{I}_{\rm out}\rvert=p_{\lvert\mathcal{I}_{\rm out}\rvert}m$ for $p_{m}\in\{0.3,0.5,0.7\}$ and $p_{\lvert\mathcal{I}_{\rm out}\rvert}\in\{0.7,0.5\}$, where we used RMSE $\lVert\widehat{\bm{X}}-\bm{X}^{\star}\rVert_{F}/\sqrt{n_{1}n_{2}}$ to measure a recovery error with the final estimate $\widehat{\bm{X}}\in\mathbb{R}^{n_{1}\times n_{2}}$ of algorithm. We note that Problem 1.1 becomes challenging as $p_{m}$ is smaller and $p_{\lvert\mathcal{I}_{\rm out}\rvert}$ is larger.
From Table 1, we observe that convergence speeds of the proposed models (5) and the model (16) are much slower than that of the model (3). This is because singular value decompositions (SVD) of $n_{1}$-by-$n_{2}$ matrices are required at every iteration in solvers for (5) (see also Remark 1.3) and for (16) while any SVD is not required in a subgradient method for (3) because of the decomposition $\bm{X}=\bm{U}\bm{V}^{\mathsf{T}}$ in (3) (Note: acceleration of the proposed Alg. 1 is beyond the scope of this paper, and will be addressed in future work). Moreover, from Table 1, the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ achieves smaller RMSE than the model (5) with $\ell\coloneqq\lvert\cdot\rvert$ in a shorter runtime; and RMSE of the existing models (3) and (16) are much higher than that of the model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ in particular for $p_{m}\in\{0.3,0.5\}$. These observations imply that a weakly convex $r^{\rm SCAD}_{\theta}$ is more robust against outliers, at least in the proposed model (5), than a convex $\lvert\cdot\rvert$ as we expected; and the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ outperforms the existing models (3) and (16) in terms of recovery accuracy [^5] We note that theoretical recovery guarantees of optimization models involving nonconvex loss functions, e.g., SCAD, have been reported for special cases of RLRMR (2), e.g., robust matrix completion [41]. However, such a theoretical recovery guarantee of the proposed model (5) is beyond the scope of this paper, and will be addressed in future work. .